OpenStudy (anonymous):

A jar contains 1 red marble and 1 marble. you grab 1 marble jar without looking. What is the sample space for this experiment?

OpenStudy (***[isuru]***):

The sample space of an experiment is the set of all possible outcomes of that experiment. the probability of getting red marble = 1/2 the probability of getting the other marble = 1/2 if you understand how those probabilities were taken.. then sample space is 1/2 + 1/2 = 1 since its the set of all possibilities.. got it :) ?

OpenStudy (anonymous):

so the answer is 1

OpenStudy (***[isuru]***):

nope srry its incomplete.. :( the set of all possibilities is 1 which means { red marble , the other marble } this is the answer sample space = { red marble , the other marble }

OpenStudy (anonymous):

thanks

OpenStudy (***[isuru]***):

welcome :)